a) Solve the following equation
0.5x+0.2(x-20)=0.3x+6
b)Solve the following equation
2/3x+1/5x=x-5/12

Answer by Garree10(2)

0.5x+0.2(x-20)=0.3x+6 times 2with -20=-4
0.5x+0.2x-4=0.3x+6
-0.2 -0.2
___________________
0.3x-4=0.3x +6